OpenStudy (anonymous):

lease find the value of c that makes the expression a perfect square trinomial, then write the expression as a square of a binomial x^2+50x+c

OpenStudy (anonymous):

x^2 + 2*25x + c For c=25^2

OpenStudy (anonymous):

To find C, you must half the middle term, then square it. So (50/2)=625. Therefore, the trinomial expression is x^2+50x+625.